# An Exploratory Case Study on Data Breach Journalism

### v1

@@ Abstract.
ccs: Security and privacy Human and societal aspects of security and privacy
@@ 2.1. Crime News and Crime Journalism
More often, however, data breaches nowadays occur online through exploitation of technical vulnerabilities in software systems. Such exploitation shifts the focus more explicitly to cyber crime. This characterization is true also in media within which big global data breaches frequently attain flashy headlines.
@@
Crime news and crime journalism have a long history, as does the fields that study this form of journalism. Already in the early 20th century (if not earlier), it was common to speculate about and argue over the role crime plays and should play in media and journalism, how crime should be reported, what function crime news fulfills or should fulfill in a society, and what consequences it may entail, among other related topics of interest. While crime journalism was an integral historical part in the popularization of media, it was also commonly seen as a hindrance to raising the status of journalism and the professionalization of journalists along the way (Rossland07, ). To this end, it was commonly argued that crime news should move away from sensationalism, which may increase the risk of emulations and copycats, among other things, toward constructive reporting that informs the people about the dangers in their neighborhoods, possibly leading to beneficial regulatory or other changes for both the public and law enforcement (Rice29, ). This characterization captures many typical tenets of crime journalism even today: Crime is newsworthy because it is, by construction, sensational, extraordinary, and disruptive; crime tells also about the political realities people live under; and crime news often requires resonance with the everyday life of people, and a way to achieve such resonance is to rely on geographical or historical proximity (Higgins22, ). It is not difficult to relate these tenets to cyber crime and data breaches in particular. Data breaches are often sensational because they showcase the banality of the cyber (in)security of many organizations; data breaches can act—and have acted—as a major political force pushing new cyber security regulations onward; and historical proximity is guaranteed because everyone in practice knows a previous breach—if not a victim of such a breach.
@@
Recently, the debate has also touched victims of crime and victimization, meaning the process of becoming a victim or being victimized by someone. Here, to begin with, exposure to sensational (or negative) crime news tends to increase the perception of people about the probability of becoming a victim of a crime (Velasquez20, ). This effect is presumably partially due to the pervasive notion of risk in crime news, which often report on serious and seemingly random crimes committed by strangers (OLeary21, ). Again, it is not difficult to draw an analogy to data breaches; if it can happen to everyone, it can happen to me as well. Though, it must be added that considerable skepticism has been expressed over whether people consume crime news because they want to learn about crime statistics and related robust details about crime or because they want to entertain themselves, protect collective moral identity, moralize on political conflicts, or something along these lines (Katz87, ). Such consumer appetite, in turn, has led to a common criticism that crime news tend to create an oversimplified or outright false picture of crime, rather promoting existing fears, stereotypes, biases, and prejudices (OLeary21, ; Smolej08, ). The use of victims (instead of criminals or suspects) in crime news may help to mitigate such perhaps unintended consequences. Regarding data breaches, it is indeed not difficult to find examples of human interest stories whereby a breach is approached through its consequences to the victims.
@@
Also other affected people may appear in crime news, including, of course, law enforcement or other public authorities, but also eyewitnesses and people related to a suspect or a victim often appear in crime news. For instance, eyewitnesses can be used in crime news to give a mediated witness experience whereby the consumers of the news align their emotional response to those affected by a crime (vanKrieken22, ). Although such journalistic techniques may in theory appear also in cyber crime and data breach news, it is more common to see a voice from law enforcement in these news. In fact, for reasons soon to be discussed, data breach news often rely on official statements from law enforcement as the building blocks for the news. Also interviews of law enforcement officials, data protection authorities, and cyber security experts sometimes appear in cyber crime and data breach news.
@@ 2.2. Potential Consequences
In light of the previous discussion, it is helpful to consider the case in a large context of data breach journalism and its potential individual, economic, societal, and other consequences. Thus, the following seven points briefly consider the potential consequences from data breach journalism in a larger context.
First, it may be that data breach journalism shares the same sins as crime news generally. In particular, it may be biased in various ways. For instance, it may exaggerate the security risks from data breaches, potentially leading to uncalled fears among the public. Consequently, data breach journalism may contribute to painting of the Internet as an excessively insecure and dangerous place. However, it is impossible to evaluate how realistic these scenarios are; the evidence is in the eye of the beholder.
@@
Third, data breach journalism may cause additional harm to the victims. Here, it is important to stress that many—if not most—data breaches involve two types of “victims”: The companies or other organizations who were breached and the persons, customers, business associates, or other types of stakeholders whose data was breached. The potential harms differ between the two types. Although the evidence is not entirely conclusive, data breaches may even affect the stock prices of the victim companies who were breached (Spanos16, ). Analogously, losing sensitive data of business associates to criminals is hardly good business. Regulations such as the General Data Protection Regulation (GDPR) have also introduced potentially large financial sanctions from data breaches. Therefore, it is no wonder that data breaches are a significant reputation threat to companies (Syed19, ). However, it is difficult or even impossible to say whether a company is still better off with a data breach compared to a ransomware attack whereby the company’s data is encrypted and kept as a ransom. Nevertheless, in both cases the potential consequences provide a clear incentive for companies to keep the information disclosed to media as limited as possible. In terms of the other victim type, a good albeit morbid case would be the Finnish Vastaamo breach. It involved a breach of a psychotherapy center whose patient records were leaked to the darknet and further used as extortion material. Given the sensitivity of the data involved and the psychological state of the victims, the constant news coverage of the breach may have caused additional stress to the victims—or patients. This case aligns with arguments on the elevated victimization of data breach victims after a notification of a breach (Gibson23, ). The example also demonstrates that data breach journalism involves its own ethical considerations.
@@
Fourth, data breach journalism may also help the user-style victims by raising awareness among them. Indeed, information about data breaches has been observed to correlate with security awareness (Chua21, ; Mayer23, ). A correlation has been further found to awareness of data protection regulations, including the GDPR (YanChua20, ). Then: If social security numbers and bank account details have been breached, a risk of identity thefts may be significant enough to warrant reporting also in media. By becoming aware of a breach, the persons affected may then take necessary precautions. Unfortunately, however, this line of reasoning does not necessarily hold that well because existing results also hint about poor practical security posture among users even after they have become aware of an incident through media or by other means (Bhagavatula21, ; Mayer23, ). Another way to look at the awareness question is to focus on the other “victims”; data breach journalism may help users also by shedding light on companies or organizations with haphazard security practices. The major Equifax breach in the United States serves as an example along these lines, which again resemble the constructive crime journalism ideals.
@@
Sixth, data breach investigations require trustworthy frameworks and secure communication channels between the parties involved, including, but not necessarily limited to, the victim organization, the investigators, and possibly law enforcement or other governmental agencies (Thomas22, ). If sensitive details leak to media during an investigation, flashy headlines may follow but these may hamper the given investigation. Such scenarios again underline that data breach journalism should follow its own ethical practices.
@@ 2.3. Related Work
According to a fairly comprehensive literature search, no directly comparable previous works exist. Having said that, there are some studies that come close to the present paper’s theme and its framing. Mimicking existing work for quantifying general crime news (Lindgren08, ), attempts have been made to profile different cyber crimes through online news coverage (Christian22, ). According to this data mining branch of research, moreover, headlines in media tend to portray negatively the victim organizations of data breaches and cyber crime in general (Mazrouei22, ). This result is hardly surprising in the sense that negative news commonly surpass positive ones, partially because negativity often drives consumption (Robertson23, ; Trussler14, ). Perhaps due to this demand-side, even many journalism textbooks tend to rationalize the commonplace reporting of crime, conflict, and bad news (Parks19, ).
@@ 4.1. Sample Characteristics
Four points about the sample are a good way to start the empirical exploration. Thus, first, Fig. 1 shows the number of news articles published annually. As can be seen, the outlet has kept a steady hpublishing cycle without notable ebbs and flows; on average, about two news articles were published each day. Regarding time series trends in general, it has been hypothesized and observed that recent events and transformations, such as COVID-19 and increased remote work, may have increased the amount of data breaches (Ozer24, ). However, such conclusions are uncalled for with the present case study; data breaches are one thing and journalistic reporting on them another. There is also a further point worth making about the longitudinal aspects. Namely, data breaches take a long time to be detected; according to some studies, even more than a hundred days is typically required before a data breach is detected within an organization (Roumani22, ). Anecdotal evidence on some major breaches support such results (Jones24, ). If no leaks occur to journalists, supposedly a long time is further required before news about a breach appear in media. Thus, the publication dates shown do not likely say much about the actual dates of the data breaches discussed in the articles.
@@
Second, as can be concluded from Fig. 2, the news articles are typically very short. In fact, the mean length, as measured by the number of tokens (that is, words), is about the same as the length of a typical abstract in a scientific paper. Although also longer news articles are present, it can be concluded that the typical articles do not have much journalistic ambition. Instead, these focus on plain reporting on data breaches, short and simple.
@@
Fourth, a few interesting observations can made based on the sources on which the news articles are based. A simple way to approximate the sources is to focus on the domain names that appear in the uniform resource locators (URLs) embedded to hyperlinks in the body of the news articles, excluding self-references to other news articles on databreaches.net. The top-15 most referenced domains are shown in Fig. 3. The most notable observation is present in the text embedded to the middle of the figure; most cases are referenced with a single unique URL. This observation hints that most of the cases are referenced with primary sources that likely point toward the victim organizations. Although such referencing clearly increases quality, there are also quite a few references to X/Twitter. It may not be a big sin, however, given that also mainstream media nowadays tend to often reference this social media platform. Otherwise, a commonplace English-language bias seems to be present, which is expected. Technology news sites in English, the public broadcaster from Canada, the data protection authority in the United Kingdom, and several governmental websites in the United States have been relatively frequently referenced in the news articles. All are quality sources. Finally, it should be also remarked that the site occasionally uses also traditional journalistic referencing, mentioning a source for a news article in plain text, without providing associated hyperlinks to the online origins.
@@ 4.2. Topics
The bigrams provide sufficient material to deduce abut the typical cases discussed in the data breach news articles. To this end, Fig. 4 displays the top-25 bigrams present in the pre-processed sample. In addition, Table 1 shows the sectors of the victim organizations according to existing meta-data provided by the news site.
Figure 4. Top-25 Bigrams
@@
  • Although information security incidents have been a common cause for GDPR fines (Ruohonen22IS, ), only seventeen news articles have contained the character string gdpr. This observation likely reflects the English language bias and the origins of the news site in the United States. Indeed, the character string hipaa, as for the Health Insurance Portability and Accountability Act (HIPAA) in the United States, already accounts for 221 mentions in the data breach news articles.
@@
The healthcare data breaches warrant a brief discussion too. Although financial crime may be involved due to stolen health insurance details, the risks associated are likely different for many victims. As was already noted, leakage of patient records may cause psychological stress to many victims. These can be used also on more serious crimes, such as extortion schemes. Another big question is why hospitals and healthcare organizations in general are so prone to data breaches and cyber attacks in general. Alas, the reasons are many. Security awareness, human competencies, organizational postures, and financial resources are problems, but these are further exacerbated by lack of oversight bodies and poor governance of healthcare systems (GarziaPerez23, ; Lee23, ), among other things. For the time being, it seems sensible to expect widespread insecurity of hospitals and healthcare facilities also in the foreseeable future. The data breaches presumably do not stop in this sector, no matter how difficult the consequences are for the victims of the breaches.

### v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: The 19th International Conference on Availability, Reliability and Security; July 30-August 2, 2024; Vienna, Austriabooktitle: The 19th International Conference on Availability, Reliability and Security (ARES 2024), July 30-August 2, 2024, Vienna, Austriadoi: 10.1145/3664476.3670456isbn: 979-8-4007-1718-5/24/07ccs: Security and privacy Human and societal aspects of security and privacy
@@ 2.1. Crime News and Crime Journalism
More often, however, data breaches nowadays occur online through exploitation of technical vulnerabilities in software systems. Such exploitation shifts the focus more explicitly to cyber crime. This characterization is true also in media within which big global data breaches frequently attain flashy headlines. Furthermore, data breaches form their own specific cyber security news niche already because these are almost always tied to specific events, unlike many other cyber threat narratives in media, such as those related to national security and espionage (Boholm21, ). In other words, a data breach in itself is almost always the specific event to which the given data breach news build upon. A data breach event typically also remains as the focal point in follow-up news.
@@
Crime news and crime journalism have a long history, as does the fields that study this form of journalism. Already in the early 20th century (if not earlier), it was common to speculate about and argue over the role crime plays and should play in media and journalism, how crime should be reported, what function crime news fulfills or should fulfill in a society, and what consequences it may entail, among other related topics of interest. While crime journalism was an integral historical part in the popularization of media, it was also commonly seen as a hindrance to raising the status of journalism and the professionalization of journalists along the way (Rossland07, ). To this end, it was commonly argued that crime news should move away from sensationalism, which may increase the risk of emulations and copycats, among other things, toward constructive reporting that informs the people about the dangers in their neighborhoods, possibly leading to beneficial regulatory or other changes for both the public and law enforcement (Rice29, ). This characterization captures many typical tenets of crime journalism even today: crime is newsworthy because it is, by construction, sensational, extraordinary, and disruptive; crime tells also about the political realities people live under; and crime news often require resonance with the everyday life of people, and a journalistic way to achieve such resonance is to rely on geographical or historical proximity (Higgins22, ). It is not difficult to relate these tenets to cyber crime and data breaches in particular. Data breaches are often sensational because they showcase the banality of the cyber (in)security of many organizations; data breaches can act—and have acted—as a major political force pushing new cyber security regulations onward; and historical proximity is guaranteed because everyone in practice knows a previous breach—if not a victim of such a breach.
@@
Recently, the debate has also touched victims of crime and victimization, meaning the process of becoming a victim or being victimized by someone. Here, to begin with, exposure to sensational (or negative) crime news tends to increase the perception of people about the probability of becoming a victim of a crime (Velasquez20, ). This effect is presumably partially due to the pervasive notion of risk in crime news, which often report on serious and seemingly random crimes committed by strangers (OLeary21, ). Again, it is not difficult to draw an analogy to data breaches; if it can happen to everyone, it can happen to me as well. Though, it must be added that considerable skepticism has been expressed over whether people consume crime news because they want to learn about crime statistics and related robust details about crime or because they want to entertain themselves, protect collective moral identity, moralize on political conflicts, or something along these lines (Katz87, ). Such consumer appetite is a central factor behind a common criticism that crime news tend to create an oversimplified or outright false picture of crime, rather promoting existing fears, stereotypes, biases, and prejudices (OLeary21, ; Smolej08, ). That said, there is also criticism of the criticism, particularly with respect to the assumption about the effect of crime news upon audiences (Doyle06, ). In any case, the use of victims (instead of criminals or suspects) in crime news may perhaps help to mitigate unintended consequences. Regarding data breaches, it is indeed not difficult to find examples of human interest stories whereby a breach is approached through its consequences to the victims.
@@
Also other affected people may appear in crime news, including, of course, law enforcement or other public authorities, but also eyewitnesses and people related to a suspect or a victim often appear in crime news. For instance, eyewitnesses can be used in crime news to give a mediated witness experience whereby the consumers of the news align their emotional response to those affected by a crime (vanKrieken22, ). Although such journalistic techniques may in theory appear also in cyber crime and data breach news, it is more common to see a voice from law enforcement in these news. In fact, for reasons soon to be discussed, data breach news often rely on official statements from law enforcement as the building blocks for the news. This sourcing may lead to a known bias, given that a law enforcement perspective is not alone the only relevant viewpoint to crime (Sacco95, ). Therefore, it is worth further remarking that not only interviews of law enforcement officials but also interviews of data protection authorities and cyber security experts sometimes appear in cyber crime and data breach news.
@@ 2.2. Potential Consequences
In light of the previous discussion, it is helpful to consider the case in a larger context of data breach journalism and its potential individual, economic, societal, and other consequences. Thus, the following seven points briefly consider the potential consequences from data breach journalism in a larger context.
First, it may be that data breach journalism shares the same sins as crime news generally. In particular, it may be biased in various ways. For instance, it may exaggerate the security risks from data breaches, potentially leading to uncalled fears among the public. Consequently, data breach journalism may contribute to painting of the Internet as an excessively insecure and dangerous place. These points align with an argument that cyber crime tends to be often associated with a moral panic in media, which may also have unintended consequences, such as inefficient resource allocation or poorly justified legal powers granted to law enforcement agencies (Lavorgna19, ). However, it is impossible to evaluate how realistic these points are in practice; the evidence is in the eye of the beholder.
@@
Third, data breach journalism may cause additional harm to the victims. Here, it is important to stress that many—if not most—data breaches involve two types of victims: the companies or other organizations who were breached and the persons, customers, business associates, or other types of stakeholders whose data was breached. The potential harms differ between the two types. Although the evidence is not entirely conclusive, data breaches may even affect the stock prices of the victim companies who were breached (Spanos16, ). Analogously, losing sensitive data of business associates to criminals is hardly good business. Regulations such as the General Data Protection Regulation (GDPR) have also introduced potentially large financial sanctions from data breaches. Therefore, it is no wonder that data breaches are a significant reputation threat to companies (Syed19, ). To protect their images, companies and other organizations typically use different crisis communication strategies when dealing with data breaches, and these may also affect the narratives presented in media (Ruohonen24PRR, ). However, it is difficult or even impossible to say whether a company is still better off with a data breach compared to a ransomware attack whereby the company’s data is encrypted and kept as a ransom. Nevertheless, in both cases the potential consequences provide a clear incentive for companies to keep the information disclosed to media as limited as possible. In terms of the other victim type, a good albeit morbid case would be the Finnish Vastaamo breach. It involved a breach of a psychotherapy center whose patient records were leaked to the darknet and further used as extortion material. Given the sensitivity of the data involved and the psychological state of the victims, the constant news coverage of the breach may have caused additional stress to the victims—or patients. This case aligns with arguments on the elevated victimization of data breach victims after a notification of a breach (Gibson23, ). The example also demonstrates that data breach journalism involves its own ethical considerations.
@@
Fourth, data breach journalism may also help the user-style victims by raising awareness among them. Indeed, information about data breaches has been observed to correlate with security awareness (Chua21, ; Mayer23, ). A correlation has been further found to awareness of data protection regulations, including the GDPR (YanChua20, ). Then: if social security numbers and bank account details have been breached, a risk of identity thefts may be significant enough to warrant reporting also in media. By becoming aware of a breach, the persons affected may then take necessary precautions. Unfortunately, however, this line of reasoning does not necessarily hold that well because existing results also hint about poor practical security posture among users even after they have become aware of an incident through media or by other means (Bhagavatula21, ; Mayer23, ). Another way to look at the awareness question is to focus on the other victims; data breach journalism may help users also by shedding light on companies or organizations with haphazard security practices. The major Equifax breach in the United States serves as an example along these lines, which again resemble the constructive crime journalism ideals.
@@
Sixth, data breach investigations require trustworthy frameworks and secure communication channels between the parties involved, including, but not necessarily limited to, the victim organization, the investigators, and possibly law enforcement or other governmental agencies (Thomas22, ). If sensitive details leak to media during an investigation, wild headlines may follow but these may hamper the given investigation. Such scenarios again underline that data breach journalism should follow its own ethical practices. Ethics apply also to the relatively frequent cases in which a data breach is found by a non-malicious third-party who may report it both to the given organization and media (IBM23, ; Ruohonen24PRR, ). In terms of ethics, the issue is rather similar to voluntary reporting of vulnerabilities to vendors. Moreover, regulations such as the GDPR have introduced notifications to the natural persons affected; therefore, leaks to media are likely to occur in any case. That said, leaks should not downplay ethics in journalism and associated professional practices.
@@ 2.3. Related Work
According to a fairly comprehensive literature search, no directly comparable previous works exist. Having said that, there are some studies that come close to the present paper’s theme and its framing. Mimicking existing work for quantifying general crime news (Lindgren08, ), attempts have been made to profile different cyber crimes through online news coverage (Christian22, ). According to this data mining branch of research, moreover, data breaches are often reported negatively in media, including with respect to the victim organizations (Mazrouei22, ; Sinanaj14, ). Such results are hardly surprising in the sense that negative news commonly surpass positive ones, partially because negativity often drives consumption (Robertson23, ; Trussler14, ). Perhaps due to this demand-side, even many journalism textbooks tend to rationalize the commonplace reporting of crime, conflict, and bad news (Parks19, ). These results reinforce the earlier point about the reputation threat from data breaches.
@@ 4.1. Sample Characteristics
Four points about the sample are a good way to start the empirical exploration. Thus, first, Fig. 1 shows the number of news articles published across the time period observed. As can be seen, the outlet has kept a steady publishing cycle; on average, about two news articles were published each day. That said, there have been some notable gaps in the publishing over the years. As a small outlet, a lack of sufficient human resources may explain these gaps.
@@
Regarding time series trends in general, it has been hypothesized and observed that recent events and transformations, such as COVID-19 and increased remote work, may have increased the amount of data breaches (Ozer24, ). However, such conclusions are uncalled for with the present case study; data breaches are one thing and journalistic reporting on them another. There is also another point worth making about the longitudinal aspects: data breaches take a long time to be detected. According to some studies and industry reports, even more than a hundred days is typically required before a data breach is detected within an organization (IBM23, ; Roumani22, ). If no leaks occur to journalists, supposedly a long time is further required before news about a breach appear in media. Thus, the publication dates shown do not likely say much about the actual dates of the data breaches discussed in the news articles.
Second, as can be concluded from Fig. 2, the news articles are typically very short. In fact, the mean length, as measured by the number of tokens (that is, words), is about the same as the length of a typical abstract in a scientific paper. Although also longer news articles are present, it can be concluded that the typical articles do not perhaps have much journalistic ambition. Instead, these focus on plain reporting on data breaches, short and simple.
@@
Figure 4. Publication Dates of Twitter References
@@
Fourth, a few interesting observations can made based on the sources on which the news articles are based. A simple way to approximate the sources is to focus on the domain names that appear in the uniform resource locators (URLs) embedded to hyperlinks in the body of the news articles, excluding self-references to other news articles on databreaches.net. The top-15 most referenced domains are shown in Fig. 3. The most notable observation is present in the text embedded to the middle of the figure; most cases are referenced with a single unique URL. This observation hints that most of the cases are referenced with primary sources that likely point toward the victim organizations. Although such referencing clearly increases quality, there are also quite a few references to X/Twitter. It may not be a big sin, however, given that also mainstream media nowadays tend to often reference this social media platform. It may also be that X/Twitter has gained prominence in data breach reporting. However, according to Fig. 4, most of the references were made in the late 2000s during which also X/Twitter’s popularity peaked. Otherwise, a commonplace English language bias seems to be present, which is expected. Technology news sites in English, the public broadcaster from Canada, the data protection authority in the United Kingdom, and several governmental websites in the United States have been relatively frequently referenced in the news articles. All are quality sources. Finally, it should be also remarked that the site occasionally uses also traditional journalistic referencing, mentioning a source for a news article in plain text, without providing associated hyperlinks to the online origins.
@@ 4.2. Topics
The bigrams provide sufficient material to deduce abut the typical cases discussed in the data breach news articles. To this end, Fig. 5 displays the top-25 bigrams present in the pre-processed sample. In addition, Table 1 shows the sectors of the victim organizations according to existing meta-data provided by the news site. None of the sectors exhibit notable longitudinal trends; the data breaches have affected the five sectors rather uniformly across time.
Figure 5. Top-30 Bigrams
@@
  • Although information security incidents have been a common cause for GDPR fines (Ruohonen22IS, ), only seventeen news articles have contained the character string gdpr. This observation likely reflects the English language bias and the origins of the news site in the United States. Indeed, the character string hipaa, as for the Health Insurance Portability and Accountability Act (HIPAA) in the United States, already accounts for 221 mentions in the data breach news articles. Also the California Consumer Privacy Act (CCPA), which was motivated by the GDPR, has gathered one reference.
@@
The healthcare data breaches warrant a brief discussion too. Although financial crime may be involved due to stolen health insurance details, the risks associated are likely different for many victims. As was already noted, leakage of patient records may cause psychological stress to many victims. These can be used also on more serious crimes, such as blackmailing schemes. In terms of the victim organizations, on the other hand, healthcare data breaches are particularly costly for the organizations involved (IBM23, ). Another big question is why hospitals and healthcare organizations in general are so prone to data breaches and cyber attacks in general. Alas, the reasons are many. Security awareness, human competencies, organizational postures, and financial resources are problems, but these are further exacerbated by lack of oversight bodies and poor governance of healthcare systems (GarziaPerez23, ; Lee23, ), among other things. For the time being, it seems sensible to expect widespread insecurity of hospitals and healthcare facilities also in the foreseeable future. The data breaches presumably do not stop in this sector, no matter how difficult the consequences are for the victims of the breaches.